OpenStudy (anonymous):

Find a formula, indexed from i = 1, for the general term ai of each of the following sequences (assuming that the pattern of the rst few terms continues).

OpenStudy (anonymous):

\[\left\{ 7,3,\frac{ 11 }{ 5},\frac{ 13 }{ 7 },\frac{ 5 }{ 3 },... \right\}\]